A 1720 W radiant heater is constructed to operate at 120 V. (a) What will be the current in the heater? (b) What is the resistance of the heating coil? (c) How much thermal energy is generated in 1.0 h?
Answer #1

Here ,

Power , P = 1720 W

V = 120 V

a) let the current in the heater is I

P = V * I

1720 = 120 * I

I = 14.33 A

the current in the heater is 14.33 A

b)

resistance = Voltage/current

resistance = 120/14.33

resistance = 8.37 Ohm

c)

thermal energy = P * time

thermal energy = 1720 * 1 * 3600

thermal energy = 6192000 J